# Simultaneous Calibration of Noise Covariance and Kinematics for State Estimation of Legged Robots via Bi-level Optimization

### v2

@@ V-B Cost Definition and Gradient Computation
Given that the full trajectory lies on the Lie-group manifold $\mathcal{M}=\mathrm{SE}(3)^{T+1}\times\mathbb{R}^{\,n_{\mathrm{eucl}}(T+1)}$ with $\mathbf{X}\in\mathcal{M}$, we denote the per-time-step state manifold by $\mathcal{G}=\mathrm{SE}(3)\times\mathbb{R}^{\,n_{\mathrm{eucl}}}$ with $\mathbf{x}\in\mathcal{G}$, which implies that $\mathcal{M}=\mathcal{G}^{T+1}$. The smooth operators on the manifold $\mathcal{G}$ are defined as $\boxplus:\ \mathcal{G}\times\mathbb{R}^{6+n_{\mathrm{eucl}}}\to\mathcal{G},\quad\boxminus:\ \mathcal{G}\times\mathcal{G}\to\mathbb{R}^{6+n_{\mathrm{eucl}}}$. We write $\log\!:\!\mathrm{SE}(3)\!\to\!\mathfrak{se}(3)$ and $\exp\!:\!\mathfrak{se}(3)\!\to\!\mathrm{SE}(3)$ for the Lie logarithm and exponential maps.
@@ V-B1 Upper level loss function $\mathcal{L}$
$=\frac{1}{2}\sum_{k=0}^{T}\left(\big\|\left(\log\big(\mathbf{T}_{k}\,\mathbf{T}_{\mathrm{GT},k}^{-1}\big)\right)^{\vee}\big\|_{2}^{2}+\big\|\mathbf{e}_{k}-\mathbf{e}_{\mathrm{GT},k}\big\|_{2}^{2}\right),$
@@ V-B2 Gradient
$\frac{{}^{e}D\mathcal{L}(\hat{\mathbf{x}})}{D\hat{\mathbf{x}}}=\lim_{\epsilon\to 0}\frac{\mathcal{L}\!\big(\hat{\mathbf{x}}\boxplus(\epsilon\tau)\big)-\mathcal{L}(\hat{\mathbf{x}})}{\epsilon},$ (15)
for any $\tau\in T_{e}(\mathcal{G})\cong\mathbb{R}^{6+n_{\mathrm{eucl}}}$. For the $\mathrm{SE}(3)$ component, with canonical basis $\{\bm{\xi}_{i}\}\subset\mathfrak{se}(3)$,
$\left[\frac{{}^{e}D\mathcal{L}(\mathbf{T})}{D\mathbf{T}}\right]_{i}=\lim_{\epsilon\to 0}\frac{\mathcal{L}\big(\exp(\epsilon\bm{\xi}_{i})\cdot\mathbf{T}\big)-\mathcal{L}(\mathbf{T})}{\epsilon}.$ (16)
@@
$\frac{{}^{e}D\hat{\mathbf{x}}(\bm{\theta})}{D\bm{\theta}}=\lim_{\epsilon\to 0}\frac{\hat{\mathbf{x}}(\bm{\theta}+\epsilon\nu)\;\boxminus\;\hat{\mathbf{x}}(\bm{\theta})}{\epsilon},$ (17)
@@
$\left[\frac{{}^{e}D\mathbf{T}}{D\bm{\theta}}\right]_{j}=\lim_{\epsilon\to 0}\frac{\left(\log\!\big(\mathbf{T}\cdot\tilde{\mathbf{T}}_{j}^{-1}\big)\right)^{\vee}}{\epsilon},$ (18)
@@
Analytical Gradient Computation: The above-mentioned limit definitions allow finite-differencing for computation. However, with long horizons and high-dimensional upper-level optimization variables, the computational cost of numerical differentiation remains considerable even when parallelized on modern CPUs. Therefore, we employ analytical derivatives instead. The upper-level loss is aggregated over time via the product-manifold residual in (14). Using the chain rule on $\mathcal{M}=\mathcal{G}^{T+1}$,
@@
$\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})\;=\;\left(\frac{D\mathcal{L}(\hat{\mathbf{x}}(\bm{\theta}))}{D\hat{\mathbf{x}}}\right)\left(\frac{D\hat{\mathbf{x}}(\bm{\theta})}{D\bm{\theta}}\right)\;+\;\frac{\partial\mathcal{L}}{\partial\bm{\theta}},$ (19)
where $D\mathcal{L}(\hat{\mathbf{x}})/D\hat{\mathbf{x}}\in\mathbb{R}^{1\times d}$ and $D\hat{\mathbf{x}}(\bm{\theta})/D\bm{\theta}\in\mathbb{R}^{d\times m}$ are formed and stacked over the horizon with $d\!=\!(6+n_{\mathrm{eucl}})(T\!+\!1)$. To differentiate through the lower-level estimator, the optimal solution $\hat{\mathbf{x}}^{\star}(\bm{\theta})$ is implicitly defined by the first-order optimality conditions of problem (8). We treat the noise variables $\mathbf{w}$, $\mathbf{n}$ and the multipliers $\bm{\lambda}$ as implicit functions of the state $\hat{\mathbf{x}}^{\star}$ and the parameter $\bm{\theta}$. The KKT residual is: $\mathcal{F}\!\left(\hat{\mathbf{x}}^{\star},\bm{\theta},\mathbf{y}(\bm{\theta}),\mathbf{G}(\bm{\theta})\right)=0$. Because of the kinematics offset parameters, the observation $\mathbf{y}$ (10) and the mapping matrices from joint kinematics to measurement space $\mathbf{G}\!=\!\{G_{i}\}$ (III-A2) depends explicitly on $\bm{\theta}_{\mathrm{foot},j}$ (13). By differentiating this system with respect to $\bm{\theta}$, we obtain the sensitivity of the optimal state trajectory by implicit function theorem:
@@
$\frac{\partial\hat{\mathbf{x}}^{\star}}{\partial\bm{\theta}}=-\;\left(\frac{\partial\mathcal{F}}{\partial\hat{\mathbf{x}}^{\star}}\right)^{-1}\left(\frac{\partial\mathcal{F}}{\partial\bm{\theta}}+\frac{\partial\mathcal{F}}{\partial\mathbf{y}}\frac{\partial\mathbf{y}}{\partial\bm{\theta}}+\frac{\partial\mathcal{F}}{\partial\mathbf{G}}\frac{\partial\mathbf{G}}{\partial\bm{\theta}}\right),$
where the right-hand side collects explicit parameter derivatives, implicit measurement dependence, and the additional contribution through $\mathbf{G}$. This states that the total effect of $\bm{\theta}$ on the optimal inner variables is the sum of a direct term and two chain–rule terms passing through $\mathbf{y}(\bm{\theta})$ and $\mathbf{G}(\bm{\theta})$.
In our estimation formulation, the decision vector $\mathbf{x}$ stacks all states over the whole window of data. Because dynamics and measurements couple locally, the KKT system is large yet sparse with a banded or block-tridiagonal pattern. Therefore, instead of forming dense matrix inverses, we express
$\left(\frac{\partial\mathcal{F}}{\partial\hat{\mathbf{x}}^{\star}}\right)\,Z=-\left(\frac{\partial\mathcal{F}}{\partial\bm{\theta}}+\frac{\partial\mathcal{F}}{\partial\mathbf{y}}\frac{\partial\mathbf{y}}{\partial\bm{\theta}}+\frac{\partial\mathcal{F}}{\partial\mathbf{G}}\frac{\partial\mathbf{G}}{\partial\bm{\theta}}\right),$
where we use sparse linear solvers to solve $Z:=(\partial\hat{\mathbf{x}}^{\star})/(\partial\bm{\theta})$ efficiently. This makes each application of the inverse a low incremental cost, achieved through back-substitutions with a high-performance solver such as PyPardiso.
@@ VI Results
We evaluate our calibration approach on three robots: a bipedal robot STRIDE [kang24] and two quadrupedal robot Go1 and B1 from Unitree. The data used from STRIDE and Go1 are obtained from simulation in Matlab and MuJoCo environment in their open source repository [kang24]. We obtained the data of B1 on hardware in the motion capture room with 12 Opti-track Cameras, with a combination of Prime 13 and 22.
The bi-level optimization is implemented in Python and C++. The outer loop iteration as LMO is solved by MOSEK; the inner loop FIE is solved with IPOPT [wachter2006implementation]. The analytical gradients are generated via CasADi [Andersson2019] and Pinocchio [carpentier:hal-01866228].
@@
Joint Sensor Noise & Kinematics Calibration: For the robot STRIDE, we mainly evaluate the joint calibration on the sensor noises and robot kinematics since their dynamics processes are deterministic in Matlab. We inject sensor noises into the reading along with an articulated shin length offset to the estimator. Fig. 4 shows the calibration results. The loss successfully reduced through the iterations of the bi-level optimization, and the estimated velocity converges to the ground truth; the kinematic parameters converge as well.
@@
Joint Noise & Kinematics Calibration: Last and most importantly, we calibrate the robot B1 using hardware data. The robot is controlled to walk in the motion-capture room. The torso motion is measured by the motion capture system, taken as ground truth with a constant unknown offset to the base frame position of the torso. The orientation measurement of the torso is assumed to be unbiased, because the torso and the marked rigid body are well-aligned in the beginning of the experiment. Additionally, we assume all the shins of B1 have kinematic errors from the factory URDF. Thus, besides the noise covariances, the kinematic errors to be calibrated are 15-dimensional. After we obtained the sensory data along with the ground truth measurements, we applied our calibration to B1. Fig. 7 shows the improved estimates of the torso’s orientation and linear velocity, and Fig. 6 shows the kinematics converge to reasonable values: in our experimental setup, the marked torso location is indeed about centimeters away from the base frame location. Table I shows the quantitative results in terms of RMSE; Fig. 8 shows the new evaluation on a different segment of data using calibrated covariances and kinematics, which still yields highly accurate estimates. To sum up, our joint calibration significantly improves the estimates by very large margins.

### v3

@@ V-B Cost Definition and Gradient Computation
Given that the full trajectory lies on the Lie-group manifold $\mathcal{M}=\mathrm{SE}(3)^{T+1}\times\mathbb{R}^{\,n_{\text{aug}}(T+1)}$ with $\mathbf{X}\in\mathcal{M}$, we denote the per-time-step state manifold by $\mathcal{G}=\mathrm{SE}(3)\times\mathbb{R}^{\,n_{\text{aug}}}$ with $\mathbf{x}\in\mathcal{G}$. The smooth operators on the manifold $\mathcal{G}$ are defined as $\boxplus:\ \mathcal{G}\times\mathbb{R}^{6+n_{\text{aug}}}\to\mathcal{G},\quad\boxminus:\ \mathcal{G}\times\mathcal{G}\to\mathbb{R}^{6+n_{\text{aug}}}$. We write $\log\!:\!\mathrm{SE}(3)\!\to\!\mathfrak{se}(3)$ and $\exp\!:\!\mathfrak{se}(3)\!\to\!\mathrm{SE}(3)$ for the Lie logarithm and exponential maps.
@@ V-B1 Upper level loss function $\mathcal{L}$
$=\frac{1}{2}\sum_{k=0}^{T}\left(\big\|\left(\log\big(\mathbf{T}_{k}\,\mathbf{T}_{\mathrm{GT},k}^{-1}\big)\right)^{\vee}\big\|_{2}^{2}+\big\|\mathbf{e}_{k}-\mathbf{e}_{\mathrm{GT},k}\big\|_{2}^{2}\right),\vskip-3.0pt$
@@ V-B2 Gradient
$\frac{{}^{e}D\mathcal{L}(\hat{\mathbf{x}})}{D\hat{\mathbf{x}}}=\lim_{\epsilon\to 0}\frac{\mathcal{L}\!\big(\hat{\mathbf{x}}\boxplus(\epsilon\tau)\big)-\mathcal{L}(\hat{\mathbf{x}})}{\epsilon},$
for any $\tau\in T_{e}(\mathcal{G})\cong\mathbb{R}^{6+n_{\text{aug}}}$. For the $\mathrm{SE}(3)$ component, with canonical basis $\{\bm{\xi}_{i}\}\subset\mathfrak{se}(3)$,
$\left[\frac{{}^{e}D\mathcal{L}(\mathbf{T})}{D\mathbf{T}}\right]_{i}=\lim_{\epsilon\to 0}\frac{\mathcal{L}\big(\exp(\epsilon\bm{\xi}_{i})\cdot\mathbf{T}\big)-\mathcal{L}(\mathbf{T})}{\epsilon},\forall i=1\cdots 6.$
@@
$\frac{{}^{e}D\hat{\mathbf{x}}(\bm{\theta})}{D\bm{\theta}}=\lim_{\epsilon\to 0}\frac{\hat{\mathbf{x}}(\bm{\theta}+\epsilon\nu)\;\boxminus\;\hat{\mathbf{x}}(\bm{\theta})}{\epsilon},$
@@
$\left[\frac{{}^{e}D\mathbf{T}}{D\bm{\theta}}\right]_{j}=\lim_{\epsilon\to 0}\frac{\left(\log\!\big(\mathbf{T}\cdot\tilde{\mathbf{T}}_{j}^{-1}\big)\right)^{\vee}}{\epsilon},\qquad\forall j=1\cdots n_{\theta},\vskip-3.0pt$
@@
Analytical Gradient Computation: The limit definitions above enable multi-threaded finite-difference approximations on modern CPUs. However, for long horizons and high-dimensional upper-level parameters, numerical differentiation is still prohibitively expensive, even with CPU parallelization. We therefore compute analytical gradients.
The upper-level objective (7) aggregates residuals on the product manifold $\mathcal{G}^{T+1}$. Applying the chain rule yields:
@@
$\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})=\left(\frac{D\mathcal{L}}{D\hat{\mathbf{x}}}\right)\left(\frac{D\hat{\mathbf{x}}}{D\bm{\theta}}\right)+\frac{\partial\mathcal{L}}{\partial\bm{\theta}},$ (8)
where the Jacobians $\frac{D\mathcal{L}}{D\hat{\mathbf{x}}}\in\mathbb{R}^{1\times r}$ and $\frac{D\hat{\mathbf{x}}}{D\bm{\theta}}\in\mathbb{R}^{r\times n_{\theta}}$ are stacked over the horizon with $r=(6+n_{\text{aug}})(T+1)$.
To differentiate through the lower-level estimator, we analyze the first-order optimality conditions of (4). Let $\mathbf{z}:=(\hat{\mathbf{x}},\mathbf{w},\mathbf{n},\bm{\lambda})$ collect the primal variables and Lagrange multipliers. We define the KKT residual mapping as
$\mathcal{F}(\mathbf{z}^{\star}(\bm{\theta}),\bm{\theta},\mathbf{y}(\bm{\theta}),\mathbf{G}(\bm{\theta}))=0.\vskip-2.0pt$ (9)
where $\mathbf{y}$ and $\mathbf{G}=\{G_{i}\}$ are the observation term (2) and the kinematics-to-measurement mapping (3), respectively. Due to the kinematic offset parameters $\bm{\theta}_{\mathrm{foot},j}$ in (6), both $\mathbf{y}$ and $\mathbf{G}$ depend explicitly on $\bm{\theta}$. Assuming $\mathcal{F}$ is continuously differentiable in a neighborhood of the solution and the Jacobian $\partial\mathcal{F}/\partial\mathbf{z}$ is nonsingular at $\mathbf{z}^{\star}(\bm{\theta})$, the implicit function theorem guarantees the existence of a locally unique differentiable solution map $\mathbf{z}^{\star}(\bm{\theta})$. Differentiating (9) with respect to $\bm{\theta}$ yields the linear system for trajectory sensitivity:
@@
$\left(\frac{\partial\mathcal{F}}{\partial\mathbf{z}}\right)\frac{\partial\mathbf{z}^{\star}}{\partial\bm{\theta}}=-\left(\frac{\partial\mathcal{F}}{\partial\bm{\theta}}+\frac{\partial\mathcal{F}}{\partial\mathbf{y}}\frac{\partial\mathbf{y}}{\partial\bm{\theta}}+\frac{\partial\mathcal{F}}{\partial\mathbf{G}}\frac{\partial\mathbf{G}}{\partial\bm{\theta}}\right),$ (10)
where $\partial\mathcal{F}/\partial\bm{\theta}$ denotes the partial derivative with $(\mathbf{y},\mathbf{G})$ held fixed. The trajectory sensitivity $\frac{D\hat{\mathbf{x}}}{D\bm{\theta}}$ is obtained by extracting the block of $\partial\mathbf{z}^{\star}/\partial\bm{\theta}$ corresponding to $\hat{\mathbf{x}}^{\star}$. In practice, (10) is solved by performing a sparse factorization of the KKT matrix $\partial\mathcal{F}/\partial\mathbf{z}$ with PyPardiso, yielding a negligible per-iteration computational cost at the upper level.
@@ VI Results
We evaluate our calibration approach on three robots: a bipedal robot STRIDE [17] and two quadrupedal robot Go1 and B1 from Unitree. The data used from STRIDE and Go1 are obtained from simulation in Matlab and MuJoCo environment in their open source repository [17]. We obtained the data of B1 on hardware in the motion capture room with 12 Opti-track Cameras, with a combination of Prime 13 and 22. The bi-level optimization is implemented in Python and C++. The outer loop iteration as LMO is solved by MOSEK; the inner loop FIE is solved with IPOPT [34]. The analytical gradients are generated via CasADi [2] and Pinocchio [8].
@@
Joint Sensor Noise & Kinematics Calibration: For the robot STRIDE, we mainly evaluate the joint calibration on the sensor noises and robot kinematics since their dynamics processes are deterministic in Matlab. We injected additive white Gaussian noise into the joint encoder angle/velocity measurements, and corrupt the IMU with non-diagonal correlated Gaussian noise, along with an articulated shin length offset to the estimator. Fig. 4 shows the calibration results. The loss successfully reduced through the iterations of the bi-level optimization, and estimated velocity converges to the ground truth; the kinematic parameters converge as well.
@@
Joint Noise & Kinematics Calibration: Last and most importantly, we calibrate the robot B1 using hardware data. The robot is controlled to walk in the motion-capture room. The torso motion is measured by the motion capture system, taken as ground truth with a constant unknown offset to the base frame position of the torso. The orientation measurement of the torso is assumed to be unbiased, because the torso and the marked rigid body are well-aligned in the beginning of the experiment. Additionally, we assume all the shins of B1 have kinematic errors from the factory URDF. Thus, besides the noise covariances, the kinematic errors to be calibrated are 15-dimensional. After we obtained the sensory data along with the ground truth measurements, we applied our calibration to B1. Fig. 7 shows the improved estimates of the torso’s orientation and linear velocity, and Fig. 6 shows the kinematics converge to reasonable values: in our experimental setup, the marked torso location is indeed about centimeters away from the base frame location. Table I shows the quantitative results in terms of RMSE; Fig. 8 shows the new evaluation on a different segment of data using calibrated covariances and kinematics, which still yields highly accurate estimates. On an Intel Core Ultra 9 275HX CPU (24C/24T) with a $15$ s window at $200$ Hz ($N=3000$ states, $n_{\theta}=69$ parameters), the proposed bi-level calibration typically converges within $\sim$15–25 outer iterations and takes about 0.5-1.5 hour wall-clock time, excluding the one-time CasADi function generation for analytical derivatives. We believe this runtime can be further reduced by exploiting the log-det barrier structure of the covariance variables and implementing a tailored interior-point solver, which may yield locally quadratic convergence once the iterates enter a well-conditioned regime. To sum up, our joint calibration significantly improves the estimates by very large margins.